# Probing of EoS with clusters and hypernuclei

### v1

@@ Abstract
The study of the nuclear equation-of-state (EoS) is a one of the primary goals of experimental and theoretical heavy-ion physics. The comparison of recent high statistics data from the STAR Collaboration with transport models provides a unique possibility to address this topic in a yet unexplored energy domain. Employing the microscopic n-body Parton-Hadron-Quantum-Molecular Dynamics (PHQMD) transport approach, which allows to describe the propagation and interactions of hadronic and partonic degrees of freedom including cluster and hyper-nucleus formation and dynamics, we investigate the influence of different EoS on bulk observables, the multiplicity, $p_{T}$ and rapidity distributions of protons, $\Lambda$s and clusters up to A=4 as well as their influence on the collective flow. We explore three different EoS: two static EoS, dubbed ’soft’ and ’hard’, which differ in the compressibility modulus, as well as a soft momentum dependent EoS. We find that a soft momentum dependent EoS reproduces most baryon and cluster observables, including the flow observables, quantitatively.
@@ I Introduction
On the theoretical side the EoS of nuclear matter can be calculated by the Brueckner-Dirac approach [9] or by chiral effective theories [10]. Both relay, however, on expansion schemes, which limit their predictive power to densities below or close to normal nuclear matter density. Lattice gauge calculations should be the tool to extend these calculations towards higher densities or chemical potentials, $\mu$, but the sign problem makes calculations difficult for $\mu\neq 0$ and at densities, which are reached in heavy ion collisions, the present efforts to determine the EoS in terms of a $\mu/T$ expansion around $\mu/T=0$ are not very reliable.
@@
In the past in heavy-ion collisions especially three observables have been identified, which are sensitive to the equation of state: subthreshold $K^{+}$ production [17, 18], the in-plane $v_{1}$[19, 20] and the elliptic flow $v_{2}$ The latter two can be obtained by expanding the azimuthal distribution of hadrons in a Fourier series
@@ II.1 QMD Propagation
[…]
where $H$ is the N-body Hamiltonian. Eq. (4) can be formally solved by approximating the N-body wave function by the direct product of single particle ”trial” wave functions (neglecting antisymmetrization) $\psi=\prod_{i}^{N}\psi_{i}$. With the assumption that the wave functions have a Gaussian form and that the width of the wave function is time independent, one obtains two equations-of-motion for the time evolution of the centroids of the Gaussian single particle Wigner density, which resemble the EoM of a classical particle with the phase space coordinates ${\bf r_{i0},p_{i0}}$ [34]:
@@
×$\times{\rm e}^{-\frac{L}{2}[\mathbf{p}_{i}^{T}(t)-\mathbf{p}_{i0}^{T}(t)]^{2}}{\rm e}^{-\frac{L}{2\gamma_{cm}^{2}}[\mathbf{p}_{i}^{L}(t)-\mathbf{p}_{i0}^{L}(t)]^{2}},$
@@
×$\times{\rm e}^{-\frac{\gamma_{cm}^{2}}{L}[\mathbf{r}_{i0}^{L}(t)-\mathbf{r}_{j0}^{L}(t)]^{2}}.$
@@ II.2 Modeling the EoS within PHQMD
Figure 1: Top: Schrödinger equivalent optical potential $U_{opt}$ versus the momentum $p$ of the proton extracted from pA collisions [50, 47, 48]. Bottom; The distribution of relative momenta among the nucleons for different times and for Au+Au collisions at $\sqrt{s_{\rm NN}}=3\rm GeV$.
@@
with $\Delta p=\sqrt{({\bf p}-{\bf p}_{1})^{2}}$. The parameterization II (blue dashed line) is the same as the parameterization I for momenta $\leq 2$ GeV$/c$, but with a constant for momentum $>2$ GeV$/c$. The parameterization III (red solid line) for $\Delta p<1.7\rm GeV$ is the same as the old one, while for $\Delta p>\rm 1.7$ GeV$/c$, we take,
@@ II.3 Relation of the potential to the EoS of nuclear matter
[…]
@@
As discussed, our equation contains the minimal number of 3 parameters $\alpha,\beta,\gamma$, which have to be determined. They are obtained by the requirement that at normal nuclear density $E/A=-16$ MeV. The compressibility modulus K is a free parameter. The hard (K = 380 MeV), the soft and the soft momentum-dependent (K=200 MeV) equations-of-state are illustrated in Fig. 2 and the parameters for the three equations-of-state are presented in Table 1. For cold nuclear matter the soft and soft momentum-dependent EoS have by construction the same $\frac{E}{A}(\rho)$.
@@ II.4 Cluster production in PHQMD
1) Coalescence mechanism: A proton and a neutron can form a deuteron if their distance at the time, when the last one of the two freezes out, is less than $|r_{1}-r_{2}|\leq 3.575$ fm in coordinate space and less than $|p_{1}-p_{2}|\leq 285$ MeV$/c$ in momentum space. For details of this cluster formation algorithm in PHQMD we refer to Ref. [31]. This method is not applied in this article.
@@
3) Potential mechanism: The attractive potential between baryons with a small relative momentum keeps them close together and can lead to a group of bound nucleons. At a given time $t$ a snapshot of the positions and momenta of all nucleons is recorded and the MST clusterization algorithm is applied: two nucleons $i$ and $j$ with the corresponding isospin are considered as “bound” to a deuteron or to a larger cluster $A>2$ if they fulfill the condition [34]
@@ III Model study: the influence of uncertainnesses in $U_{opt}(p)$ on observables
Figure 4: The in-plane flow $v_{1}(p_{T})$ near target rapidity ($-1<y<-0.5$) of protons and deuterons from PHQMD simulations of Au+Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV (left column) and 5.4 GeV (right column) in the 10–40% centrality class. Results are shown for three different parameterizations of the nuclear optical potential (Parameterization I, II, and III).
Figure 5: The elliptic flow $v_{2}(p_{T})$ at mid-rapidity ($-0.5<y<0$) of protons and deuterons from PHQMD simulations of Au+Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV (left column) and 5.4 GeV (right column) in the 10–40% centrality class. Results are shown for three different parameterizations of the nuclear optical potential (Parameterization I, II, and III).
@@
As discussed in Section II.B there are no experimental data, which would allow for a reliable extrapolation of $U_{opt}(p)$ to large $p$, see Fig. 1. However, a momentum $p>2$ GeV$/c$ is achievable already at the beam energies of $\sqrt{s}$ = 3 GeV - as follows from the lower plot in Fig. 1. In our model study we explore the 3 parametrizations of $U_{opt}(p)$, shown in Fig. 1, and investigate whether the different parametrization lead to different values of the observables for heavy-ion reactions at $\sqrt{s_{\rm{NN}}}=3$ GeV as well and at 5 GeV, which allow to explore even larger $p$.
@@ IV Centrality selection within PHQMD
To compare data and model calculations, we determine the centrality in PHQMD in the same way as done in the experimental analysis. In experimental data, the centrality class is defined using the measured charged particle multiplicity (FXTMult) within the acceptance of the STAR Time Projection Chamber (TPC) ($-2.0<\eta<0$). The relationship between FXTMult and collision centrality is determined by fitting Glauber model calculations [62] to FXTMult. In PHQMD, charged particles ($\pi,~{}K,~{}p$) are selected within the same $\eta$ range as in data, and with 0.2 $<p_{\rm T}<$ 10.0 GeV/c and $p_{\rm T}>$ 0.35 GeV/c for protons, which mimics the acceptance of the STAR detector. Subsequently, the PHQMD multiplicity distribution is scaled to match the $\sqrt{s_{\rm{NN}}}=3$ GeV STAR data, as shown in Fig. 6. As can be seen, the distribution for the different EoS is rather similar to each other and to the data. The tail in the data at FXTMult$>200$ is due to pileup events, i.e. multiple collision events, which occur within a short period of time or within the same detector readout window. These events are not included in the experimental analysis. The FXTMult values, corresponding to different collision centralities, are determined using the following formula:
[…]
@@
Figure 6: The multiplicity distribution of charged particles in Au+Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV of STAR data (black dots) and of the hadronic transport model PHQMD using a hard (green band), a soft (blue band) and a soft momentum dependent (red) equation of state (EoS). Vertical lines indicate the minimum number of tracks required for an event to be in the corresponding centrality bin for the soft EoS.
@@ V.1 Transverse Momentum Spectra
We start out with the comparison of measured differential yields of various particles to PHQMD calculations for several centrality bins with different EoS. Figures 7 and  8 show the $p_{T}$ spectra of proton ($p$), deuteron ($d$), triton ($t$), ${}^{3}\rm{He}$, ${}^{4}\rm{He}$, and $\Lambda$. For protons, calculations using H and SM EoS describe the data fairly well. The agreement between data and theory increases with increasing centrality and approaching midrapidity. Very good agreement is found near mid-rapidiy, and in central collisions. The predictions underestimate the data near target rapidity in more peripheral collisions. There the emission of light clusters happens by fragmentation or by thermal emission on a much longer time scale than the passing of the two nuclei in heavy ion reactions. The time scale for cluster production at the rapidity of the residue is given the time to regain the spatial form of the residue and by deexcitation of excited residue states. This time scale is much longer than the time we can pursue the simulations. Below $p_{T}=1$ GeV$/c$ calculations using a S EoS are similar to those using a H and SM EoS. For $p_{T}>1$ GeV$/c$, the spectra with a S EoS are softer than those with a H and SM EoS, and the S EoS calculations underestimate the data in all rapidity and centrality bins. For deuteron, triton, ${}^{3}\rm{He}$, and ${}^{4}\rm{He}$, the overall trends are similar. However, deuterons are better described by the model compared to heavier nuclei (A $\geq$ 3). For ${}^{3}\rm{He}$, and ${}^{4}\rm{He}$, the calculations systematically underpredict the data across rapidity and centrality bins. The comparison of the measured $\Lambda$ transverse momentum spectra to the model calculations is, in general, quite similar to those of protons. Calculations using S EoS underestimate the data above $p_{T}=1$ GeV$/c$ except for the most peripheral bin. The calculations using H and SM EoS slightly overestimates the data in central collisions at mid-rapidity in central collisions.
@@
Figure 10: The coalescence parameters $B_{A}$ to the power $(A-1)^{-1}$ for deuterons (upper panel), tritons (middle panel) and ${}^{3}\rm{He}$ (bottom panel) as a function of the $p_{T}$ divided by $A$ in $0-10\%$ Au+Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV. The data are shown as black markers and the blast-wave fit to the data is shown as dash-dotted lines. The calculations from PHQMD using soft, hard and soft with momentum dependence EoS are shown as blue, green and red lines respectively.
@@ V.2 $p_{T}$-Integrated Yields And Yield Ratios
In the experimental data analysis, the $p_{T}$-integrated yields are estimated by extrapolating the $p_{T}$ spectra to the unmeasured region using functions like the blast-wave formula [63],
@@
[…]
where $m_{T}$ is the transverse mass of the particle, $I_{0}$ and $K_{1}$ are the modified Bessel functions, and $\rho(r)=tanh^{-1}(\beta_{T})$, where $\beta_{T}$ is the radial flow velocity. In the region $0\leq r\leq R$, $\beta_{T}$ can be expressed as $\beta_{T}=\beta_{S}(r/R)^{n}$, where $\beta_{S}$ is the surface velocity, and $n$ reflects the form of the flow velocity profile. For the experimental data reported in this manuscript, $n$ is fixed to 1, i.e. the transverse velocity profile is chosen to depend linearly on the radius, while $T_{kin}$, $\beta_{T}$, and the overall normalization are free parameters in the fit. The fitted values for various particles observed in the $0-10\%$ centrality bin at mid-rapidity $y=(-0.1,0)$ are tabulated in Tab. 2.
Particles $T_{\text{kin}}$ (GeV) $\langle\beta\rangle$ (c)
$p$ $0.062\pm 0.004$ $0.45\pm 0.01$
$\Lambda$ $0.062\pm 0.006$ $0.39\pm 0.02$
$d$ $0.070\pm 0.004$ $0.40\pm 0.01$
$t$ $0.078\pm 0.005$ $0.37\pm 0.01$
3He $0.077\pm 0.005$ $0.38\pm 0.01$
4He $0.080\pm 0.010$ $0.36\pm 0.02$
Table 2: Fitted kinetic freeze-out parameters for various particles, measured in $0-10\%$ centrality bin at mid-rapidity $y=(-0.1,0)$
for $\sqrt{s_{\rm{NN}}}=3$ GeV Au+Au collisions.
@@
is used instead, where $m_{T}=\sqrt{m_{0}^{2}+p_{T}^{2}}$ is the transverse mass of the particle, and $T_{m_{T}}$ is a fit parameter. Based on the difference of the estimated yields using different extrapolation functions, a systematic uncertainty is assigned. It depends both on the statistical accuracy of the $p_{T}$ spectra and the fraction of the unmeasured region. For the majority of cases, this uncertainty ranges from $10-30\%$.
Fig. 11 shows the $p_{T}$-integrated $p$, $d$, $t$, ${}^{3}\rm{He}$, ${}^{4}\rm{He}$, $\Lambda$, ${}^{3}_{\Lambda}\rm{H}$ and ${}^{4}_{\Lambda}\rm{H}$ yields as a function of rapidity for $0-10\%$ central Au+Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV. As in the previous section, the data [13, 11, 15] are compared to PHQMD with a S, H, and SM EoS. It should be noted that for PHQMD, we directly use the yield without any extrapolation. Thus, differences of the rapidity distribution between PHQMD calculations and the data can arise from two factors: 1) a difference of the integrated yield in the measured region 2) a difference of the extrapolation of the experimental data into the unmeasured $p_{T}$ regions by a fit function because the extrapolation function is not uniquely determined. The latter may be a consequence of a different shape of the calculations and data in the measured region, even if both agree in between the error bars.
@@
Figure 13: The yield ratios of hypernuclei to nuclei as a function of the mass number in $0-10\%$ Au+Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV [13, 15, 11]. The measured data points are shown as black markers, while the calculations from PHQMD using soft, hard, and soft with momentum dependence EoS are shown as blue, green, and red bands respectively. Expectations from the thermal model [64] are also shown.
Figure 14: The yield ratios $N_{p}\times N_{t}/N_{d}^{2}$ (left) and $N_{{}^{4}\rm{He}}\times N_{p}/(N_{{}^{3}\rm{He}}\times N_{d})$ (right) in the rapidity region $y=(-0.5,0)$ as a function of centrality in Au+Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV [11]. The calculations from PHQMD using soft, hard, and soft momentum dependent EOS are shown as blue, green, and red bands, respectively. Calculations from transport models with coalescence afterburner [65] and thermal model [64] are also shown for comparison.

### v2

@@ Abstract
[…]
@@ I Introduction
On the theoretical side the EoS of nuclear matter can be calculated by the Brueckner-Dirac approach [9] or by chiral effective theories [10]. Both rely, however, on expansion schemes, which limit their predictive power to densities below or close to normal nuclear matter density. Lattice gauge calculations should be the tool to extend these calculations towards higher densities or chemical potentials, $\mu$, but the sign problem makes calculations difficult for $\mu\neq 0$ and at densities, which are reached in heavy ion collisions, the present efforts to determine the EoS in terms of a $\mu/T$ expansion around $\mu/T=0$ are not very reliable.
@@
In the past in heavy-ion collisions especially three observables have been identified, which are sensitive to the equation of state: subthreshold $K^{+}$ production [17, 18], the in-plane $v_{1}$[19, 20] and the elliptic flow $v_{2}$ . The latter two can be obtained by expanding the azimuthal distribution of hadrons in a Fourier series
@@ II.1 QMD Propagation
$\color[rgb]{0,0,0}\delta\int_{t_{1}}^{t_{2}}dt\langle\psi(t)|i\frac{d}{dt}-H|\psi(t)\rangle=0$}, (4)
where $H$ is the N-body Hamiltonian and $\langle\ldots\rangle$ means the expectation value with respect to the many-body wavefunction (same as the next equation). Eq. (4) can be formally solved by approximating the N-body wave function by the direct product of single particle ”trial” wave functions (neglecting antisymmetrization) $\psi=\prod_{i}^{N}\psi_{i}$. With the assumption that the wave functions have a Gaussian form and that the width of the wave function is time independent, one obtains two equations-of-motion for the time evolution of the centroids of the Gaussian single particle Wigner density, which resemble the EoM of a classical particle with the phase space coordinates $\bf r_{i0},p_{i0}$} [34]:
@@
$\times{\rm e}^{-\frac{L}{2}[\mathbf{p}_{i}^{T}(t)-\mathbf{p}_{i0}^{T}(t)]^{2}}{\rm e}^{-\frac{L}{2\gamma_{cm}^{2}}[\mathbf{p}_{i}^{L}(t)-\mathbf{p}_{i0}^{L}(t)]^{2}},$
@@
$\times{\rm e}^{-\frac{\gamma_{cm}^{2}}{L}[\mathbf{r}_{i0}^{L}(t)-\mathbf{r}_{j0}^{L}(t)]^{2}}.$
@@ II.2 Modeling the EoS within PHQMD
Figure 1: Top: Schrödinger equivalent optical potential $U_{opt}$ versus the absolute value of relative momentum $p$ of the proton extracted from pA collisions [50, 47, 48]. Bottom; The distribution of the relative momenta among the nucleons for different times and for Au+Au collisions at $\sqrt{s_{\rm NN}}=3\rm GeV$.
@@
with $\Delta p=\sqrt{({\bf p}-{\bf p}_{1})^{2}}$. The parameterization II (blue dashed line) is the same as the parameterization I for momenta $\leq 2$ GeV/c/c, but with a constant for momentum >2>2 GeV/c/c. The parameterization III (red solid line) for $\Delta p<1.7\rm GeV$ is the same as the old one, while for $\Delta p>\rm 1.7$ GeV/c/c, we take,
@@ II.3 Relation of the potential to the EoS of nuclear matter
To calculate the energy per nucleon, we introduce $U=\int V(\rho)d\rho$. This allows to write
@@
As discussed, our equation contains the minimal number of 3 parameters $\alpha,\beta,\gamma$, which have to be determined. They are obtained by the requirement that at normal nuclear density E/A=16E/A=-16 MeV. The compressibility modulus K is a free parameter. The hard (K = 380 MeV), the soft and the soft momentum-dependent (K=200 MeV) equations-of-state are illustrated in Fig. 2 and the parameters for the three equations-of-state are presented in Table 1. For cold nuclear matter the soft and soft momentum-dependent EoS have by construction the same $\frac{E}{A}(\rho)$.
@@ II.4 Cluster production in PHQMD
1) Coalescence mechanism: A proton and a neutron can form a deuteron if their distance at the time, when the last one of the two freezes out, is less than $|r_{1}-r_{2}|\leq 3.575$ fm in coordinate space and less than $|p_{1}-p_{2}|\leq 285$ MeV/c/c in momentum space. For details of this cluster formation algorithm in PHQMD we refer to Ref. [31]. This method is not applied in this article.
@@
3) Potential mechanism: The attractive potential between baryons with a small relative momentum keeps them close together and can lead to a group of bound nucleons. At a given time $t$ a snapshot of the positions and momenta of all nucleons is recorded and the MST (Minimum Spanning Tree) clusterization algorithm is applied: two nucleons $i$ and $j$ with the corresponding isospin are considered as “bound” to a deuteron or to a larger cluster $A>2$ if they fulfill the condition [34]
@@ III Model study: the influence of uncertainnesses in $U_{opt}(p)$ on observables
Figure 4: The in-plane flow $v_{1}(p_{T})$ near target rapidity (1<y<0.5-1<y<-0.5) of protons and deuterons from PHQMD simulations of Au+Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV (left column) and 5.4 GeV (right column) in the 10–40% centrality class. Results are shown for three different parameterizations of the nuclear optical potential (Parameterization I, II, and III).
Figure 5: The elliptic flow $v_{2}(p_{T})$ at mid-rapidity (0.5<y<0-0.5<y<0) of protons and deuterons from PHQMD simulations of Au+Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV (left column) and 5.4 GeV (right column) in the 10–40% centrality class. Results are shown for three different parameterizations of the nuclear optical potential (Parameterization I, II, and III).
@@
As discussed in Section II.B there are no experimental data, which would allow for a reliable extrapolation of $U_{opt}(p)$ to large $p$, see Fig. 1. However, a momentum $p>2$ GeV/c/c is achievable already at the beam energies of $\sqrt{s}$ = 3 GeV - as follows from the lower plot in Fig. 1. In our model study we explore the 3 parametrizations of $U_{opt}(p)$, shown in Fig. 1, and investigate whether the different parametrization lead to different values of the observables for heavy-ion reactions at $\sqrt{s_{\rm{NN}}}=3$ GeV as well and at 5 GeV, which allow to explore even larger $p$.
@@ IV Centrality selection within PHQMD
To compare data and model calculations, we determine the centrality in PHQMD in the same way as done in the experimental analysis. In experimental data, the centrality class is defined using the measured charged particle multiplicity (FXTMult) within the acceptance of the STAR Time Projection Chamber (TPC) ($2.0<\eta<0$). The relationship between FXTMult and collision centrality is determined by fitting Glauber model calculations [62] to FXTMult. In PHQMD, charged particles ($\pi,~K,~p$) are selected within the same $\eta$ range as in data, and with 0.2 $p_{\rm T}<$ 10.0 GeV/c and $p_{\rm T}>$ 0.35 GeV/c for protons, which mimics the acceptance of the STAR detector. Subsequently, the PHQMD multiplicity distribution is scaled to match the $\sqrt{s_{\rm{NN}}}=3$ GeV STAR data, as shown in Fig. 6. As can be seen, the distribution for the different EoS is rather similar to each other and to the data. The tail in the data at FXTMult>200>200 is due to pileup events, i.e. multiple collision events, which occur within a short period of time or within the same detector readout window. These events are not included in the experimental analysis. The FXTMult values, corresponding to different collision centralities, are determined using the following formula:
$C_{M}=\frac{1}{\sigma^{AA}}\int^{\infty}_{M}dM^{\prime}\frac{d\sigma}{dM^{\prime}}$ (27)
@@
Figure 6: The multiplicity distribution of charged particles in Au+Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV of STAR data (black dots) and of the hadronic transport model PHQMD using a hard (green band), a soft (blue band) and a soft momentum dependent (red) equation of state (EoS). Vertical lines indicate the minimum number of tracks required for an event to be in the corresponding centrality bin for the soft EoS with momentum dependence.
@@ V.1 Transverse Momentum Spectra
Figure 7: The transverse-momentum spectra of proton for different rapidities and centralities in Au+Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV. The proton spectra have been corrected for the feed-down from $\Lambda$ decays. The measured data points from STAR [11] are shown as colored markers and the calculations from PHQMD using soft EOS, hard EOS, and soft EOS with momentum dependence are shown as dotted, dashed, and solid lines respectively. The data points and model calculations are scaled by factors of 1/10 from mid- to forward rapidities as indicated in the legend.
We start out with the comparison of measured differential yields of various particles to PHQMD calculations for several centrality bins with different EoS. Figures 7 show the proton $p_{T}$ spectra at for different rapidities and centralities in Au+Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV. The PHQMD calculations using H and SM EoS describe the data fairly well. The agreement between data and theory increases with increasing centrality and approaching midrapidity. Very good agreement is found near mid-rapidiy, and in central collisions. The predictions underestimate the data near target rapidity in more peripheral collisions. There the emission of light clusters happens by fragmentation or by thermal emission on a much longer time scale than the passing of the two nuclei in heavy ion reactions. The time scale for cluster production at the rapidity of the residue is given the time to regain the spatial form of the residue and by deexcitation of excited residue states. This time scale is much longer than the time we can pursue the simulations. Below $p_{T}=1$ GeV/c/c calculations using a S EoS are similar to those using a H and SM EoS. For $p_{T}>1$ GeV/c/c, the spectra with a S EoS are softer than those with a H and SM EoS, and the S EoS calculations underestimate the data in all rapidity and centrality bins. The deuteron, triton, $^{3}\rm{He}$, $^{4}\rm{He}$ and $\Lambda$ are shown in Appendix Fig. 25, Fig. 26, respectively, and exhibit similar overall trends.
@@
Figure 9: The coalescence parameters $B_{A}$ to the power $(A-1)^{-1}$ for deuterons (upper panel), tritons (middle panel) and $^{3}\rm{He}$ (bottom panel) as a function of the $p_{T}$ divided by AA in $0-10\%$ Au+Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV. The data are shown as black markers and the blast-wave fit to the data is shown as dash-dotted lines. The calculations from PHQMD using soft, hard and soft with momentum dependence EoS are shown as blue, green and red lines respectively.
@@ V.2 $p_{T}$-Integrated Yields And Yield Ratios
In the experimental data analysis, due to detector acceptance and statistics, the low- and high-$p_{T}$ regions cannot be measured directly. The $p_{T}$-integrated yields are therefore estimated by extrapolating the spectra to the unmeasured regions using functions such as the blast-wave formula [63],
@@
$\int_{0}^{R}rdrm_{T}$ $I_{0}(\frac{p_{T}sinh\rho(r)}{T_{kin}})K_{1}(\frac{m_{T}cosh\rho(r)}{T_{kin}})$ (30)
where $m_{T}$ is the transverse mass of the particle, $I_{0}$ and $K_{1}$ are the modified Bessel functions, and $\rho(r)=tanh^{-1}(\beta_{T})$, where $\beta_{T}$ is the radial flow velocity. In the region $0\leq r\leq R$, $\beta_{T}$ can be expressed as $\beta_{T}=\beta_{S}(r/R)^{n}$, where $\beta_{S}$ is the surface velocity, and $n$ reflects the form of the flow velocity profile. The average transverse flow velocity of the source is then defined as $\langle\beta_{T}\rangle=\frac{2}{n+2}\beta_{S}$, which represents the spatially averaged transverse velocity of the fireball over the radial profile at kinetic freeze-out. For the experimental data reported in this manuscript, $n$ is fixed to 1, i.e. the transverse velocity profile is chosen to depend linearly on the radius, while $T_{kin}$ and the overall normalization are free parameters in the fit. The fitted values for various particles observed in the $0-10\%$ centrality bin at mid-rapidity y=(0.1,0)y=(-0.1,0) are tabulated in Tab. 2, together with the fractions of the yield within the measured $p_{T}$ range relative to the total yield [11].
Particles $T_{\text{kin}}$ (GeV) $\langle\beta_{T}\rangle$ (c) Measured percentage (%)
$p$ $0.062\pm 0.004$ $0.45\pm 0.01$ $55\pm 4$
$\Lambda$ $0.062\pm 0.006$ $0.39\pm 0.02$ $55\pm 1$
$d$ $0.070\pm 0.004$ $0.40\pm 0.01$ $36\pm 5$
$t$ $0.078\pm 0.005$ $0.37\pm 0.01$ $23\pm 2$
3He $0.077\pm 0.005$ $0.38\pm 0.01$ $26\pm 2$
4He $0.080\pm 0.010$ $0.36\pm 0.02$ $15\pm 1$
Table 2: Fitted kinetic freeze-out parameters for various particles and the percentage (%) of the measured $p_{T}$ range relative to the total $dN/dy$, measured in $0-10\%$ centrality bin at mid-rapidity y=(0.1,0)y=(-0.1,0) for $\sqrt{s_{\rm{NN}}}=3$ GeV Au+Au collisions.
@@
is used instead, where $m_{T}=\sqrt{m_{0}^{2}+p_{T}^{2}}$ is the transverse mass of the particle, and $T_{m_{T}}$ is a fit parameter. Based on the difference of the estimated yields using different extrapolation functions, a systematic uncertainty is assigned. It depends both on the statistical accuracy of the $p_{T}$ spectra and the fraction of the unmeasured region. For the majority of cases, this uncertainty ranges from $3-6\%$ for protons and increases for $^{4}_{\Lambda}\rm{H}$ with $10-30\%$.
Fig. 10 shows the $p_{T}$-integrated $p$, $d$, $t$, $^{3}\rm{He}$, $^{4}\rm{He}$, $\Lambda$, $^{3}_{\Lambda}\rm{H}$ and $^{4}_{\Lambda}\rm{H}$ yields as a function of rapidity for $0-10\%$ central Au+Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV. As in the previous section, the data [13, 11, 15] are compared to PHQMD with a S, H, and SM EoS. It should be noted that in PHQMD, yields are obtained over the full $p_{T}$ range, so no extrapolation is needed. Thus, differences of the rapidity distribution between PHQMD calculations and the data can arise from two factors: 1) a difference of the integrated yield in the measured region 2) a difference of the extrapolation of the experimental data into the unmeasured $p_{T}$ regions by a fit function because the extrapolation function is not uniquely determined. The latter may be a consequence of a different shape of the calculations and data in the measured region, even if both agree in between the error bars.
@@
Figure 12: The yield ratios of hypernuclei to nuclei as a function of the mass number in $0-10\%$ Au+Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV [13, 15, 11]. The measured data points are shown as black markers, while the calculations from PHQMD using soft, hard, and soft with momentum dependence EoS are shown as blue, green, and red bands respectively. Expectations from the Thermal-Fist model [64] are also shown.
@@
Figure 13: The yield ratios $N_{p}\times N_{t}/N_{d}^{2}$ (left) and $N_{{}^{4}\rm{He}}\times N_{p}/(N_{{}^{3}\rm{He}}\times N_{d})$ (right) in the rapidity region y=(0.5,0)y=(-0.5,0) as a function of centrality in Au+Au collisions at $\sqrt{s_{\rm{NN}}}=3$ GeV [11]. The calculations from PHQMD using soft, hard, and soft momentum dependent EOS are shown as blue, green, and red bands, respectively. Calculations from transport models with coalescence afterburner [65] and thermal model [64] are also shown for comparison.